Suppose you borrow $100 today at an interest rate of 10 percent per year (paid at the end of each year). You put it into an investment that will give benefits of $60 per year for the next two years’ time (at the end of each year). In two years’ time, what will be the net benefit of your investment (the benefits minus the total costs, including interest costs).
1
Expert's answer
2020-11-17T12:17:17-0500
"2\times60-2\times 0.1\times 10=100"

Profit will be $100